# Hybrid Training for Enhanced Multi-task Generalization in Multi-agent Reinforcement Learning

### v1

@@ Introduction
We conduct experiments using the task sets with two different data qualities, expert and medium. We train online, offline, and our hybrid methods with offline data only from three source tasks and online environments respectively and evaluate them in a wide range of unseen tasks. To ensure comparable fairness, we train each method with the same 35k training step number. Detailed hyper-parameters of the experiments are available in Appendix B. The average test win rates of the marine-hard task set are shown in Table 1. The tables record the best test win rates between BC-t and BC-r as BC-best. We find that HyGen consistently outperforms other baselines in both source tasks and unseen tasks. HyGen can discover general skills from multi-task data and reuse them with high-level policies, resulting in superior and stable performance compared with UPDeT-l and UPDeT-m, which cannot generalize well among different levels of tasks. We find that BC methods and ODIS sometimes present comparable performance to HyGen, particularly with expert datasets. However, in real-world scenarios where non-expert data quality is more common, these purely offline methods are hampered by data quality limitations. This results in less robust performance and weaker cross-task generalization compared to HyGen, as clearly demonstrated by the training outcomes on the medium dataset. The results of the experiment for the stalker-zealot task set can be found in Appendix C.
Figure 4: Comparison of HyGen, QMIX, and ODIS on the source task $3m$. Both HyGen and ODIS were trained using medium-quality offline datasets. We run HyGen and ODIS for 35k steps due to early convergence and train QMIX for 100k steps. All experiment results were conducted over five random seeds.
@@
We also compare the efficiency of HyGen with the online MARL method QMIX and the offline MARL method ODIS. As illustrated in Figure 4, HyGen’s learning speed surpasses that of QMIX, demonstrating greater sample efficiency. Initially, HyGen and ODIS exhibit comparable learning speeds, but after 8k steps, HyGen becomes more efficient than ODIS. We believe that this improvement can be attributed to the increasing significance of online interactions, which continually enhance performance over time, unlike purely offline MARL methods that eventually encounter dataset limitations.
Ablation Study
@@
In our ablation studies, we investigate the effectiveness of components in our proposed HyGen structure. First, we try to find whether the linearly decreasing hybrid ratio scheme can yield better performance than the fixed. We perform HyGen hybrid training separately with the dynamic hybrid ratio and three fixed hybrid ratios which are 20%, 50%, and 80% in the marine-hard task set with medium-quality offline datasets. As we see in Figure 5, HyGen with a linearly decreasing hybrid ratio outperforms those with fixed hybrid ratios in both effectiveness and efficiency. This improvement indicates that a linearly decreasing hybrid ratio better utilizes the initial efficiency of offline learning, as offline data typically contain more useful experiences than early-stage online interactions. Starting with a higher percentage of samples from offline datasets enhances sample efficiency. Furthermore, a linearly decreasing ratio gradually increases the proportion of samples from the online replay buffer over time, progressively integrating the diversity of online interactions for exploration.
@@
We also conduct experiments to investigate our proposed action decoder refinement during the hybrid training phase. We run experiments with two variants of HyGen, with and without action decoder refinement, during the hybrid training phase in the marine-hard task set using medium-quality offline datasets and present average test win rates in source tasks and unseen tasks. Figure 6 shows that HyGen with action decoder refinement performs better, suggesting that refining skills during hybrid training reduces their dependency on the quality of the offline dataset. This improvement occurs because online interactions contribute abundant experiences that foster exploration and enhance skill diversity.
(a) Source Task: $3m$
(b) Unseen Task: $10m$
@@
We then evaluate the effectiveness of the dynamic CQL loss scheme during the hybrid training phase. We conduct experiments with HyGen using the dynamic CQL loss scheme, comparing its performance against both the fixed CQL loss and no CQL loss schemes in the marine-hard task set with medium-quality offline datasets. The superior results of HyGen with the dynamic CQL loss indicate that this approach not only mitigates the OOD problem but also avoids the impact of excessive Q-value penalties on training performance in hybrid training, as shown in Figure 7.
(a) Source Task: $3m$
(b) Unseen Task: $10m$
@@
The number of general skills is a key hyper-parameter of HyGen which we recommend to set equal to the number of attention heads according to the self-attention mechanism. To substantiate this, we conducted experiments comparing performance metrics using a fixed number of attention heads with varying skill numbers in the marine-hard task set with medium-quality offline datasets. Results indicate that equalizing skill and attention head numbers yield optimal performance, as detailed in Appendix C.
Conclusion
@@
We introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. By utilizing datasets of limited offline experiences and engaging in small-scale interactive environments, HyGen effectively discovers general skills. This approach enables the learning of a general policy applicable across diverse tasks, leading to enhanced performance in both familiar source tasks and novel, unseen tasks. Our experimental results demonstrate that HyGen effectively addresses the performance limitations inherent in offline MARL algorithms and significantly outperforms traditional online MARL algorithms in terms of efficiency. We believe that the success of HyGen underscores the importance of integrating skill discovery with hybrid training methodologies to achieve generalization in cooperative MARL scenarios and will be instrumental for the practical application of MARL in real-world settings. Future work will focus on exploring integrating large language models (LLM) to utilize their universal knowledge to further enhance the efficiency and adaptability of HyGen, particularly focusing on scalability across even more diverse and dynamic real-world applications.
Appendix A A. Descriptions of Task Sets and Offline Multi-task Datasets
Task Sets
@@
Table 2: Descriptions of tasks in the marine-hard task set. [60]
@@ 5 Zealots
Table 3: Descriptions of tasks in the stalker-zealot task set. [60]
Offline Multi-task Datasets
@@
Considering our focus on generalizing to unseen tasks, we employ offline datasets exclusively from the source tasks in the three aforementioned task sets. The Properties of offline datasets with different qualities are detailed in Table 4. Data from various tasks are amalgamated into a multi-task dataset, facilitating simultaneous multi-task policy training.
@@
Appendix B B. Experiments Details
The specific hyper-parameters of HyGen are listed in Table 5. All the tabular results show the performance of HyGen with 50, 000 optimization steps, and the steps of the hybrid high-level policy learning phase are the subtraction of the general skill discovery steps from the total steps. Our experiments are conducted on a server equipped with one Intel Xeon E5 CPU@3.60GHz processor (6 cores, 12 threads), 128 GB memory, and 2 RTX 3090 GPU cards and it usually costs 10-14 hours typically. Our HyGen code follows Apache License 2.0, the same as the PyMARL framework.
@@
Table 5: Hyper-parameters of HyGen.
Appendix C C. Additional Experiments
Experiments with different skill numbers
@@
Experiments with the stalker-zealot task set
Task Expert Medium
BC-best UPDeT-I UPDeT-m ODIS HyGen (ours) BC-best UPDeT-I UPDeT-m ODIS HyGen (ours)
Source Tasks
2s3z 93.1 $\pm$ 4 53.1 $\pm$ 39 50.0 $\pm$ 33 97.7 $\pm$ 2 97.1 $\pm$ 3 48.8 $\pm$ 9 30.6 $\pm$ 12 35.0 $\pm$ 23 49.2 $\pm$ 8 73.5 $\pm$ 11
2s4z 78.1 $\pm$ 8 48.4 $\pm$ 24 23.4 $\pm$ 26 60.9 $\pm$ 6 86.2 $\pm$ 10 12.5 $\pm$ 8 28.8 $\pm$ 4 18.8 $\pm$ 10 32.8 $\pm$ 12 51.3 $\pm$ 8
3s5z 92.5 $\pm$ 4 40.6 $\pm$ 11 17.2 $\pm$ 19 87.5 $\pm$ 9 88.9 $\pm$ 13 24.4 $\pm$ 12 15.0 $\pm$ 10 25.6 $\pm$ 24 28.9 $\pm$ 6 52.6 $\pm$ 13
Unseen Tasks
1s3z 45.6 $\pm$ 23 26.6 $\pm$ 25 1.6 $\pm$ 1 76.6 $\pm$ 3 84.1 $\pm$ 5 21.9 $\pm$ 37 33.1 $\pm$ 18 3.8 $\pm$ 5 41.4 $\pm$ 18 54.2 $\pm$ 8
1s4z 60.0 $\pm$ 32 37.5 $\pm$ 31 26.6 $\pm$ 19 17.2 $\pm$ 10 44.5 $\pm$ 9 6.2 $\pm$ 7 35.0 $\pm$ 7 2.5 $\pm$ 3 50.7 $\pm$ 7 67.3 $\pm$ 7
1s5z 45.6 $\pm$ 26 14.8 $\pm$ 13 29.7 $\pm$ 26 2.5 $\pm$ 2 47.2 $\pm$ 13 3.1 $\pm$ 2 13.1 $\pm$ 11 5.0 $\pm$ 4 14.1 $\pm$ 8 34.2 $\pm$ 13
2s5z 75.6 $\pm$ 11 27.3 $\pm$ 19 23.4 $\pm$ 22 27.3 $\pm$ 6 72.4 $\pm$ 15 14.4 $\pm$ 9 17.5 $\pm$ 9 16.9 $\pm$ 14 32.0 $\pm$ 4 43.7 $\pm$ 5
3s3z 80.6 $\pm$ 9 49.2 $\pm$ 25 20.3 $\pm$ 10 89.1 $\pm$ 5 93.3 $\pm$ 6 45.6 $\pm$ 14 23.8 $\pm$ 6 24.4 $\pm$ 28 23.4 $\pm$ 9 41.3 $\pm$ 7
3s4z 92.5 $\pm$ 5 59.4 $\pm$ 16 12.5 $\pm$ 19 96.9 $\pm$ 2 93.9 $\pm$ 5 40.0 $\pm$ 19 17.5 $\pm$ 10 28.8 $\pm$ 31 50.8 $\pm$ 15 71.6 $\pm$ 9
4s3z 67.5 $\pm$ 19 50.8 $\pm$ 24 6.2 $\pm$ 4 64.1 $\pm$ 13 74.9 $\pm$ 13 28.8 $\pm$ 26 3.1 $\pm$ 4 11.2 $\pm$ 18 13.3 $\pm$ 7 52.6 $\pm$ 13
4s4z 53.1 $\pm$ 18 41.4 $\pm$ 16 7.8 $\pm$ 13 79.7 $\pm$ 10 74.1 $\pm$ 16 20.0 $\pm$ 12 1.9 $\pm$ 2 1.2 $\pm$ 1 12.5 $\pm$ 7 44.2 $\pm$ 15
4s5z 40.6 $\pm$ 19 28.1 $\pm$ 17 5.5 $\pm$ 7 86.7 $\pm$ 12 89.9 $\pm$ 7 14.4 $\pm$ 8 5.0 $\pm$ 5 5.6 $\pm$ 8 7.0 $\pm$ 4 28.1 $\pm$ 13
4s6z 48.1 $\pm$ 23 10.9 $\pm$ 7 4.7 $\pm$ 6 88.3 $\pm$ 8 86.1 $\pm$ 12 3.8 $\pm$ 3 2.5 $\pm$ 2 1.9 $\pm$ 2 1.6 $\pm$ 1 23.8 $\pm$ 14
Table 7: Average test win rates of the final policies in the task set stalker-zealot with different data qualities, averaged over five random seeds. Results of BC-best represent the highest test win rates between BC-t and BC-r. For ease of reference, asymmetric task names are abbreviated, with ’2s3z’ denoting the SMAC map ’2s_vs_3z’.
Table 7 records the average test win rates of the stalker-zealot task set. Table records the best test win rates between BC-t and BC-r as BC-best. We find that HyGen consistently outperforms other baselines in both source tasks and unseen tasks. HyGen can discover general skills from multi-task data and reuse them with high-level policies, resulting in superior and stable performance compared with UPDeT-l and UPDeT-m, which cannot generalize well among different levels of tasks. We find that BC methods and ODIS sometimes present comparable performance to HyGen, particularly with expert datasets. However, in real-world scenarios where non-expert data quality is more common, these purely offline methods are hampered by data quality limitations. This results in less robust performance and weaker cross-task generalization compared to HyGen, as clearly demonstrated by the training outcomes on the medium dataset.

### v2

@@ 5.1.2 Experimental Results
We conduct experiments using the task sets with two different data qualities: expert and medium. We train baselines and HyGen with offline data only from three source tasks and online environments respectively and evaluate them in a wide range of unseen tasks. To ensure fairness, we train each method with the same 35k training steps. Detailed hyper-parameters of the experiments are available in Appendix B. The average test win rates of the marine-hard task set and the stalker-zealot task set are shown in Table 1. We find that HyGen generally outperforms other baselines in both source tasks and unseen tasks. HyGen can discover general skills from multi-task data and reuse them with high-level policies, resulting in superior and stable performance compared with UPDeT methods, which cannot generalize well among different levels of tasks. We find that BC methods, ODIS, and HiSSD sometimes present comparable performance to HyGen, particularly with expert datasets. However, in real-world scenarios where non-expert data quality is more common, these purely offline methods are hampered by data quality limitations. This results in less robust performance and weaker cross-task generalization compared to HyGen, as clearly demonstrated by the training outcomes on the medium dataset. In addition, HyGen significantly outperforms all other baselines in the stalker-zealot task set, a more complex task set than the marine-hard one due to the diversity of agents involved. We believe that this underscores HyGen’s superior performance on more intricate maps.
Figure 4: Comparison of HyGen, QMIX, and ODIS on the source task $3m$.
@@
We also compare the efficiency of HyGen with the online MARL method QMIX and the offline MARL method ODIS. As illustrated in Figure 4, HyGen’s learning speed surpasses QMIX’s, demonstrating greater sample efficiency. Initially, HyGen and ODIS exhibit comparable learning speeds, but after 8k steps, HyGen becomes more efficient than ODIS. We believe this improvement can be attributed to the increasing significance of online interactions, which continually enhance performance over time, unlike purely offline MARL methods that eventually encounter dataset limitations.
5.2 Ablation Study
Table 2: Ablation studies on HyGen. We report average test win rates of the final policies in the task set marine-hard with medium data qualities, averaged over five random seeds. For ease of reference, asymmetric task names are abbreviated, with ’9m10m’ denoting the SMAC map ’9m_vs_10m’.
Task 20% Fixed 50% Fixed 80% Fixed w/o Refinement w/o CQL Fixed CQL HyGen
Source Tasks
3m 63.3 $\pm$ 14 78.4 $\pm$ 8 60.4 $\pm$ 17 72.7 $\pm$ 12 56.8 $\pm$ 9 61.1 $\pm$ 19 91.5 $\pm$ 11
5m6m 20.4 $\pm$ 12 27.2 $\pm$ 18 23.9 $\pm$ 15 26.1 $\pm$ 7 21.2 $\pm$ 8 31.7 $\pm$ 9 36.1 $\pm$ 7
9m10m 45.7 $\pm$ 11 41.6 $\pm$ 12 62.5 $\pm$ 11 55.5 $\pm$ 16 46.4 $\pm$ 9 54.3 $\pm$ 17 79.2 $\pm$ 4
Unseen Tasks
4m 72.1 $\pm$ 8 79.0 $\pm$ 11 65.3 $\pm$ 10 69.9 $\pm$ 15 75.8 $\pm$ 4 82.7 $\pm$ 9 91.4 $\pm$ 8
5m 80.2 $\pm$ 11 67.1 $\pm$ 25 69.4 $\pm$ 12 53.6 $\pm$ 21 88.5 $\pm$ 7 76.9 $\pm$ 14 96.5 $\pm$ 6
10m 42.6 $\pm$ 7 76.2 $\pm$ 11 44.3 $\pm$ 9 71.1 $\pm$ 17 70.9 $\pm$ 14 47.9 $\pm$ 23 96.4 $\pm$ 3
12m 64.3 $\pm$ 12 69.6 $\pm$ 18 58.9 $\pm$ 11 55.3 $\pm$ 12 61.0 $\pm$ 7 67.3 $\pm$ 13 81.5 $\pm$ 14
7m8m 8.8 $\pm$ 7 19.8 $\pm$ 5 5.3 $\pm$ 15 5.3 $\pm$ 9 11.5 $\pm$ 8 16.5 $\pm$ 11 24.5 $\pm$ 9
8m9m 5.3 $\pm$ 11 9.6 $\pm$ 13 17.2 $\pm$ 6 15.9 $\pm$ 10 19.5 $\pm$ 11 17.1 $\pm$ 9 22.3 $\pm$ 10
10m11m 41.4 $\pm$ 14 45.8 $\pm$ 7 38.8 $\pm$ 12 26.8 $\pm$ 21 29.5 $\pm$ 18 33.1 $\pm$ 8 47.2 $\pm$ 13
10m12m 1.3 $\pm$ 4 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 2.7 $\pm$ 6 5.2 $\pm$ 2
13m15m 0.0 $\pm$ 0 5.7 $\pm$ 8 2.3 $\pm$ 2 0.0 $\pm$ 0 2.0 $\pm$ 3 3.1 $\pm$ 7 9.3 $\pm$ 6
@@ 5.2 Ablation Study
In our ablation studies, we investigate the effectiveness of components in our proposed HyGen structure. First, we try to find whether the linearly decreasing hybrid ratio scheme can yield better performance than the fixed. We perform HyGen hybrid training separately with the dynamic hybrid ratio and three fixed hybrid ratios of 20%, 50%, and 80% in the marine-hard task set with medium-quality offline datasets. As we see in Table 2 and Figure 5, HyGen with a linearly decreasing hybrid ratio outperforms those with fixed hybrid ratios in both effectiveness and efficiency. This improvement indicates that a linearly decreasing hybrid ratio better utilizes the initial efficiency of offline learning, as offline data typically contain more useful experiences than early-stage online interactions. Starting with a higher percentage of samples from offline datasets enhances sample efficiency. Furthermore, a linearly decreasing ratio gradually increases the proportion of samples from the online replay buffer over time, progressively integrating the diversity of online interactions for exploration.
@@
We also conduct experiments to investigate our proposed action decoder refinement during the hybrid training phase. We run experiments with two variants of HyGen, with and without action decoder refinement, during the hybrid training phase and present average test win rates in source tasks and unseen tasks. According to the results in Table 2, HyGen with action decoder refinement performs better, suggesting that refining skills during hybrid training reduces their dependency on the quality of the offline dataset. This improvement occurs because online interactions contribute abundant experiences that foster exploration and enhance skill diversity.
@@
We then evaluate the effectiveness of the dynamic CQL loss scheme during the hybrid training phase. We conduct experiments with HyGen using the dynamic CQL loss scheme, comparing its performance against both the fixed CQL loss and no CQL loss schemes. The superior results of HyGen with the dynamic CQL loss indicate that this approach not only mitigates the OOD problem but also avoids the impact of excessive Q-value penalties on training performance in hybrid training, as shown in Table 2.
@@
The number of general skills is a key hyper-parameter of HyGen which we recommend to set equal to the number of attention heads according to the self-attention mechanism. To substantiate this, we conducted experiments comparing performance metrics using a fixed number of attention heads with varying skill numbers in the marine-hard task set with medium-quality offline datasets. Results indicate that equalizing skill and attention head numbers yields optimal performance, as detailed in Appendix C.
6 Conclusion
@@
We introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. By utilizing datasets of limited offline experiences and engaging in small-scale interactive environments, HyGen effectively discovers general skills. This approach enables the learning of a general policy applicable across diverse tasks, leading to enhanced performance in both familiar source tasks and novel, unseen tasks. Our experimental results demonstrate that HyGen effectively addresses the performance limitations inherent in offline MARL algorithms and significantly outperforms traditional online MARL algorithms in terms of efficiency. We believe that the success of HyGen underscores the importance of integrating skill discovery with hybrid training methodologies to achieve generalization in cooperative MARL scenarios and will be instrumental for the practical application of MARL in real-world settings. Future work will focus on exploring integrating large language models (LLMs) to utilize their universal knowledge to further enhance the efficiency and adaptability of HyGen, particularly focusing on scalability across even more diverse and dynamic real-world applications.
Appendix A Descriptions of Task Sets and Offline Multi-task Datasets
A.1 Task Sets
@@
Table 3: Descriptions of tasks in the marine-hard task set. [39]
@@
Table 4: Descriptions of tasks in the stalker-zealot task set. [39]
@@ 5 Zealots
A.2 Offline Multi-task Datasets
@@
Considering our focus on generalizing to unseen tasks, we employ offline datasets exclusively from the source tasks in the three aforementioned task sets. The Properties of offline datasets with different qualities are detailed in Table 5. Data from various tasks is amalgamated into a multi-task dataset, facilitating simultaneous multi-task policy training.
@@
Appendix B Experiments Details
The specific hyper-parameters of HyGen are listed in Table 6. All the tabular results show the performance of HyGen with 50, 000 optimization steps, and the steps of the hybrid high-level policy learning phase are the subtraction of the general skill discovery steps from the total steps. Our experiments are conducted on a server equipped with one Intel Xeon E5 CPU@3.60GHz processor (6 cores, 12 threads), 128 GB memory, and 2 RTX 3090 GPU cards, and it usually costs 10-14 hours. Our HyGen code follows Apache License 2.0, the same as the PyMARL framework.
Table 6: Hyper-parameters of HyGen.
@@ Appendix B Experiments Details
Appendix C Additional Experiments with different skill numbers
@@
Appendix D Additional Results on Ablation Study
(a) Source Task: $3m$
(b) Unseen Task: $10m$
(a) Source Task: $3m$
(b) Unseen Task: $10m$